Let X be the number of applicants who apply for a senior level position at a large multinational corporation. The probabilityplease fill in the blanks?
Answer #1

Solution:-

Outcome (Number of Applicant)

0

1

2

3

4

Probability distribution

0.35

0.30

0.12

0.15

0.08

Cumulative Probability distribution

0.35

(0.35 + 0.30) =0.65

(0.35+0.30+0.12) = 0.77

(0.35+0.30+0.12+0.15) = 0.92

(0.35+0.30+0.12+0.15+0.08) = 1

Probability that there will be at least two applicants = P (2) + P (3) + P (4)

                                                                                 = 0.12 + 0.15 + 0.08

                                                                                 = 0.35

Probability that there will be at max three applicants = P (0) + P (1) + P (2) + P (3)

                                                                                   = 0.35 + 0.30 + 0.12 + 0.15

                                                                                   = 0.92

Probability that there will be at three or four applicants = P (3) + P (4)

                                                                                = 0.15 + 0.08

                                                                               = 0.23